# Towards Effectively Detecting and Explaining Vulnerabilities Using Large Language Models

### v1

@@ Abstract.
Software vulnerabilities pose significant risks to the security and integrity of software systems. Prior studies have proposed a series of approaches to vulnerability detection using deep learning or pre-trained models. However, there is still a lack of vulnerability’s detailed explanation for understanding apart from detecting its occurrence. Recently, large language models (LLMs) have shown a remarkable capability in the comprehension of complicated context and content generation, which bring opportunities for the detection and explanation of vulnerabilities of LLMs. In this paper, we conduct a comprehensive study to investigate the capabilities of LLMs in detecting and explaining vulnerabilities and propose LLMVulExp, a framework that utilizes LLMs for vulnerability detection and explanation. Under specialized fine-tuning for vulnerability explanation, our LLMVulExp not only detects the types of vulnerabilities in the code but also analyzes the code context to generate the cause, location, and repair suggestions for these vulnerabilities. We find that LLMVulExp can effectively enable the LLMs to perform vulnerability detection (e.g., over 90% F1 score on SeVC dataset) and explanation. We also explore the potential of using advanced strategies such as Chain-of-Thought (CoT) to guide the LLMs concentrating on vulnerability-prone code and achieve promising results.
Co-first authors, equally contributed.
Corresponding author.
copyright: noneconference: ; June 2024; Online
1. Introduction
@@
A software vulnerability is a flaw or weakness in a system that can be exploited by an attacker to perform unauthorized actions (Zhou et al., 2019; Li et al., 2021a). These vulnerabilities can lead to severe consequences, including data breaches, financial losses, and damage to an organization’s reputation. The increasing complexity and interconnectedness of software systems introduce the great challenge of identifying and mitigating these vulnerabilities effectively.
@@
Current vulnerability detection techniques mainly include pattern based and deep learning based approaches. Pattern based approaches (Yamaguchi, 2017; rou, [n. d.]) generally rely on manually defined rules to detect vulnerabilities. Deep learning based approaches (Zhou et al., 2019; Li et al., 2021a; Dam et al., 2017; Li et al., 2018; Russell et al., 2018) train the models using existing vulnerability data and various code representation techniques. Despite these advancements, existing methods often fall short of providing detailed explanations of detected vulnerabilities. This lack of robust explanatory capabilities impedes a comprehensive understanding and effective mitigation of vulnerabilities when applied to real-world usage. It is important to propose new techniques that can detect software vulnerabilities and provide additional explanations. Figure 1 shows an example of the vulnerability detection result with and without an explanation which provides comprehensive information for understanding and fixing the vulnerabilities.
@@
In the context of enhancing detection methods, the advent of Large Language Models (LLMs) offers a promising avenue. LLMs, with their advanced generative capabilities, have demonstrated potential in a wide range of applications, such as natural language processing and machine translation. These models can generate extensive textual content and provide contextually relevant information, making them natural candidates for tasks requiring detailed explanations. However, a substantial gap exists between the capabilities of LLMs and the specific requirements of vulnerability detection and explanation. They struggle to effectively detect vulnerabilities and provide accurate explanations without the domain knowledge of vulnerabilities, which requires a deep understanding of code structures, security contexts, and the intricate interplay between various software components. To bridge this gap, it is imperative to fine-tune LLMs specifically for vulnerability detection. This specialized training aims to endow LLMs with the necessary skills to accurately identify and explain software vulnerabilities.
@@
Figure 1. Example of vulnerability detection result with and without explanation.
@@
The challenges of this fine-tuning process are twofold: obtaining high-quality training data (Zhou et al., 2019; Chen et al., 2023; Zhou et al., 2024; Zhao et al., 2023; Wang et al., 2024) and effectively adapting the models (Xu et al., 2023; Hu et al., 2021; Thapa et al., 2022). Curating relevant datasets that encompass diverse and representative software vulnerabilities is crucial. These datasets must capture various vulnerability types, from simple coding errors to complex logic flaws. Furthermore, fine-tuning requires sophisticated techniques to tailor LLMs to the nuanced requirements of vulnerability detection and explanation. Effective fine-tuning can enhance the model’s understanding of code structures and security contexts, thereby improving its detection and explanatory capabilities.
@@
In this study, we explore enhancing LLMs to detect and explain software vulnerabilities by proposing an automated framework LLMVulExp, which involves leveraging prompt-based techniques to annotate explanatory information for open-source vulnerability data using proprietary LLM models. The annotated explanatory information includes the location of the vulnerability and a detailed explanation. Following this, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2021) fine-tuning methods to refine the LLMs, improving their detection capabilities and establishing a framework for evaluating their explanatory performance (Fu et al., 2023a). Additionally, we investigate the correlation between vulnerability detection and explanation, as well as utilizing LLMs to follow a Chain-of-Thought (CoT) (Wei et al., 2022) strategy and perform a concentrated vulnerability detection on the key code that is prone to software vulnerabilities. Furthermore, we address the evaluation of generated vulnerability explanations (Cheshkov et al., 2023) by proposing new evaluation metrics and an automated review method using LLMs.
@@
Through a comprehensive evaluation of two widely used vulnerability datasets (i.e., SeVC (Li et al., 2021c) and DiverseVul (Chen et al., 2023)), we find that LLMVulExp can enable the LLMs to detect vulnerabilities with a high accuracy (e.g., with an F1 score over 90% on SeVC) and provide an effective explanation. We also find that the key code extraction following a CoT strategy can improve the performance of vulnerability detection by a large margin.
@@
Paper Organization. Section 2 summarizes the related work. Section 3 presents the methodology of our study. Section 4 discusses the results to our research questions. Section 5 discusses the implications of our study. Section 6 discusses the threats to validity. Section 7 concludes the paper.
2. Related Work
@@
2.1. Vulnerability Detection
Prior studies proposed a series of deep learning approaches (Zhou et al., 2019; Li et al., 2021a; Dam et al., 2017; Li et al., 2018; Russell et al., 2018) to detect the vulnerabilities. These studies utilized labeled data with and without vulnerability to train neural networks and capture their semantic characteristics.
Apart from deep learning approaches, recent advancements in large language models (LLMs) have significantly influenced the field of software vulnerability detection. Advanced methodologies such as zero-shot prompting (Zhou et al., 2024; Fu et al., 2023b), in-context learning (Dong et al., 2022; Brown et al., 2020a), and fine-tuning (Hu et al., 2021; Xu et al., 2023; Chen et al., 2023) have been employed in vulnerability detection. Cheshkov et al. (Cheshkov et al., 2023) evaluated the performance of the ChaptGPT and GPT-3 on vulnerability detection tasks and found they failed to classify vulnerable codes in binary and multi-label settings. Gao et al. (Gao et al., 2023) proposed a benchmark for using LLMs in vulnerability detection and validated that on simple datasets with few-shot prompting, LLMs can achieve performance that is comparable to or exceeds that of deep learning methods. Nong et al. (Nong et al., 2024) focused on evaluating the performance improvement of vulnerability detection tasks using chain-of-thought prompting. By introducing chain-of-thought forms based on the semantic structure of code, LLMs can achieve higher detection accuracy. Sun et al. (Sun et al., 2024) further assessed the true reasoning capabilities of LLMs by decoupling their vulnerability reasoning abilities. They found that supplementing LLMs with high-quality vulnerability-related knowledge and contextual information can enhance their performance. Yusuf et al. (Yusuf and Jiang, 2024) discovered through experiments that natural language instructions enhance the performance of vulnerability detection tasks across multiple programming languages. Steenhoek et al. (Steenhoek et al., 2024) surveyed and evaluated eleven LLMs for their capabilities of vulnerability detection, utilizing various types of prompts, including in-context learning and chain-of-thought, which indicates the accuracy limitation of directly applying LLMs to vulnerability detection without finetuning.
@@ 2.2. Vulnerability Explanation
The VulDeeLocator  (Li et al., 2021b) enhances a Bi-LSTM detector by incorporating an inner multiplication layer, which aids in forecasting vulnerable statements based on the outputs of this layer. In a similar vein, both IVDetect (Li et al., 2021a) and LineVul (Fu and Tantithamthavorn, 2022) develop vulnerability detection models that utilize subgraphs or attention weights derived from the trained detectors to identify vulnerable statements. VELVET (Ding et al., 2022) prioritizes vulnerable statements by integrating graph-based neural networks with sequence-based neural networks. LineVD (Hin et al., 2022) addresses statement-level vulnerability detection as a node classification challenge, employing graph neural networks combined with a transformer-based model on PDG for comprehensive learning. VulTeller (Zhang et al., 2023) emphasizes control flows and taint flows to detect more accurate dependencies for localizing vulnerabilities. VulExplainer  (Cheng et al., 2024) is proposed to locate the fine-grained information pertaining to the vulnerability based on graph neural networks. Coca (Cao et al., 2024) is a dual-perspective contrastive learning enhancement strategy to improve vulnerability explanation methods based on graph neural networks.
@@
3. Methodology
Figure 2. Overview of our study.
To address the current gap in generative vulnerability explanation models and enhance the ability of LLMs to identify and analyze software vulnerabilities, we propose a comprehensive framework for fine-tuning and evaluating specialized models for vulnerability detection and explanation. Figure 2 presents an overview of our framework, namely, LLMVulExp. Specifically, our framework consists of four core stages: ❶ open-source vulnerability data collection, ❷ automated vulnerability explanation annotation based on prompt engineering, ❸ specialized fine-tuning of vulnerability detection and explanation through instruction-based fine-tuning, and ❹ evaluation of generative vulnerability explanation capabilities.
❶ Fine-tuning and Evaluation Vulnerability Dataset Collection: Enhancing the specialized capabilities of LLMs relies on large quantities of high-quality domain-specific data. In the context of vulnerability detection and explanation, the authenticity of the vulnerability code, the diversity of vulnerability types, and the sufficiency of examples of each type are particularly important. In this paper, we conduct the study on two datasets: (1) SeVC (Li et al., 2021c), which contains four core types and over 50,000 vulnerable code snippets, and (2) DiverseVul (Chen et al., 2023), which covers 295 real open-source projects and 150 CWE types.
Table 1. Statistics of the studied datasets.
@@ 3. Methodology
Semantics-based Vulnerability Candidate (SeVC) dataset includes 126 distinct types of vulnerabilities with 56,395 vulnerable samples and 364,232 non-vulnerable ones. The SeVC dataset is categorized into four primary groups based on the underlying causes of the vulnerabilities: Library/API Function Call, Array Usage, Pointer Usage and Arithmetic Expression. The SeVC dataset includes a significant number of vulnerabilities for each of its four types, making it suitable for fine-tuning and evaluating models designed to detect and explain specific types of vulnerabilities.
@@
DiverseVul is a C/C++ vulnerable source code dataset, which includes 18,945 vulnerable functions and 330,492 non-vulnerable functions derived from 7,514 commits, encompassing 150 Common Weakness Enumerations (CWEs). DiverseVul is currently the largest real-world C/C++ vulnerability dataset, characterized by longer code snippets, coverage of more diverse projects and CWE types, and lower label noise. Therefore, it is used to evaluate the capabilities of LLMs in handling more challenging real-world scenarios involving a broader range of vulnerability types. We select the top ten CWE types to construct our fine-tuning dataset, enabling us to conduct multi-label classification detection
@@
We deduplicate the vulnerability samples using a hash method. Then, we downsample the non-vulnerability samples in a 1:1 ratio to the vulnerability samples to obtain a balanced dataset. The dataset details are shown in Table 1. We split the processed dataset into training, validation, and test sets in an 80%: 10%: 10% ratio for both vulnerability and non-vulnerability samples to conduct our experiments.
@@
❷ Automated Vulnerability Explanation Annotation: Current open-source vulnerability datasets primarily include information within the scope of source code, vulnerability labels, CWE types, and commit messages. However, there is still a lack of detailed explanations of the vulnerability logic within the source code, which poses challenges for vulnerability detection techniques to provide a corresponding explanation regarding the detection results. Manually annotating real-world vulnerable code explanations requires extensive software development experience and in-depth knowledge of software vulnerabilities, involving high labor and time costs.
@@
To address this challenge, we propose an automated vulnerability explanation annotation method based on prompt engineering using open-source LLMs. This method leverages the contextual learning and instruction-following capabilities of LLMs, using prompt engineering to achieve large-scale, high-quality automated vulnerability explanation annotation. Prompts decompose the explanation goal into three sub-goals: vulnerability discrimination, code localization, and specific explanation. Combining instruction-based prompt templates and effective annotated examples to stimulate the model’s contextual learning capabilities ensures the effectiveness of vulnerability explanation annotation.
@@
In this paper, we use GPT-3.5 (Brown et al., 2020b) by invoking the API provided by OpenAI (ope, 2024) to implement the annotation process. To address our research questions and experimental needs, we annotated 40,491 and 9,161 vulnerability explanation data points across two datasets, respectively. This effort fills the current gap in vulnerability explanation data. The specific annotation process can be found in Section 4.1.
@@
❸ Specialized Fine-Tuning of Vulnerability Detection and Explanation: The automated explanation annotation of open-source vulnerability data provides a large-scale dataset for vulnerability detection and explanation, addressing the data bottleneck issues in the fine-tuning process. To enhance the vulnerability detection and explanation capabilities of LLMs (especially open-source models with lower computational overhead), we fine-tune general-purpose LLMs to enable them to detect and explain specific types of vulnerabilities in real code. We use instruction-based prompts to guide tasks, helping LLMs correctly understand task goals and generate standardized output. To reduce the computational overhead of the fine-tuning process, we adopt the parameter-efficient fine-tuning technique LoRA (Hu et al., 2021), significantly reducing time and space costs.
@@
❹ Evaluation of Generative Vulnerability Explanation: The current lack of research on model-generated vulnerability explanations makes effective evaluation of LLM-generated vulnerability explanations an important issue. Similar to the challenges faced in annotation, manual evaluation also requires significant human and time resources. Furthermore, the effectiveness of vulnerability explanations involves multiple dimensions to provide useful assistance to developers in real development scenarios. To efficiently evaluate vulnerability explanations, we propose an evaluation method based on three dimensions: accuracy, clarity, and constructiveness, combined with prompt engineering to develop an automated LLM evaluation method. Through expert manual verification, we validate the effectiveness of the specialized vulnerability explanation model’s generated quality and the feasibility of the LLM-based automated evaluation scheme.
3.1. Vulnerability Interpretation Enhancement Prompting

### v2

# mystyle backgroundcolor=mygray, leftmargin=0.45em, rightmargin=0.45em, innertopmargin=10pt, innerbottommargin=10pt, skipabove=10pt, skipbelow=10pt

@@ Abstract
Software vulnerabilities pose significant risks to the security and integrity of software systems. Prior studies have proposed various approaches to vulnerability detection using deep learning or pre-trained models. However, there is still a lack of detailed explanations for understanding vulnerabilities beyond merely detecting their occurrence, which fails to truly help software developers understand and remediate the issues. Recently, large language models (LLMs) have demonstrated remarkable capabilities in comprehending complex contexts and generating content, presenting new opportunities for both detecting and explaining software vulnerabilities. In this paper, we conduct a comprehensive study to investigate the capabilities of LLMs in both detecting and explaining vulnerabilities, and we propose LLMVulExp, a framework that utilizes LLMs for these tasks. Under specialized fine-tuning for vulnerability explanation, our LLMVulExp not only detects the types of vulnerabilities in the code but also analyzes the code context to generate the cause, location, and repair suggestions for these vulnerabilities. These detailed explanations are crucial for helping developers quickly analyze and locate vulnerability issues, providing essential guidance and reference for effective remediation. We find that LLMVulExp can effectively enable the LLMs to perform vulnerability detection (e.g., achieving over a 90% F1 score on the SeVC dataset) and provide detailed explanations. We also explore the potential of using advanced strategies such as Chain-of-Thought (CoT) to guide the LLMs in concentrating on vulnerability-prone code, achieving promising results.
11footnotetext:  Co-first authors, equally contributed.22footnotetext:  Corresponding author.
I Introduction
@@
A software vulnerability is a flaw or weakness in a system that can be exploited by an attacker to perform unauthorized actions [1, 2, 3, 4, 5]. These vulnerabilities can lead to severe consequences, including data breaches, financial losses, and damage to an organization’s reputation. The increasing complexity and interconnectedness of software systems introduce the great challenge of identifying and mitigating these vulnerabilities effectively.
@@
Current vulnerability detection techniques mainly include pattern based and deep learning based approaches. Pattern based approaches [6, 7] generally rely on manually defined rules to detect vulnerabilities. Deep learning based approaches [1, 2, 8, 9, 10] train the models using existing vulnerability data and various code representation techniques. Despite these advancements, existing methods often fall short of providing detailed explanations of detected vulnerabilities. This lack of robust explanatory capabilities impedes a comprehensive understanding and effective mitigation of vulnerabilities when applied to real-world usage. It is important to propose new techniques that can detect software vulnerabilities and provide additional explanations. Figure 1 shows an example of the vulnerability detection result with and without an explanation which provides comprehensive information for understanding and fixing the vulnerabilities.
@@
In the context of enhancing detection methods, the advent of Large Language Models (LLMs) offers a promising avenue. LLMs, with their advanced generative capabilities, have demonstrated potential in a wide range of applications, such as natural language processing and machine translation. These models can generate extensive textual content and provide contextually relevant information, making them natural candidates for tasks requiring detailed explanations. However, a substantial gap exists between the capabilities of LLMs and the specific requirements of vulnerability detection and explanation. They struggle to effectively detect vulnerabilities and provide accurate explanations without the domain knowledge of vulnerabilities, which requires a deep understanding of code structures, security contexts, and the intricate interplay between various software components. To bridge this gap, it is imperative to fine-tune LLMs specifically for vulnerability detection and explanation. This specialized training aims to endow LLMs with the necessary skills to accurately identify and explain software vulnerabilities.
@@
Figure 1: Example of vulnerability detection result with and without explanation.
@@
The challenges of this fine-tuning process are twofold: obtaining high-quality training data [1, 11, 12, 13, 14] and effectively adapting the models [15, 16, 17]. Curating relevant datasets that encompass diverse and representative software vulnerabilities is crucial. These datasets must capture various vulnerability types, from simple coding errors to complex logic flaws. Furthermore, fine-tuning requires sophisticated techniques to tailor LLMs to the nuanced requirements of vulnerability detection and explanation. Effective fine-tuning can enhance the model’s understanding of code structures and security contexts, thereby improving its detection and explanatory capabilities.
@@
In this study, we explore the enhancement of LLMs for detecting and explaining software vulnerabilities by proposing an automated framework LLMVulExp, which involves leveraging prompt-based techniques to annotate explanatory information for open-source vulnerability data using advanced LLM models. The annotated explanatory information includes the location of the vulnerability and a detailed explanation. Following this, we employ Low-Rank Adaptation (LoRA) [16] fine-tuning methods to refine the LLMs, improving their detection capabilities and establishing a framework for evaluating their explanatory performance [18]. Additionally, we investigate the correlation between the capability of vulnerability detection and explanation, as well as utilize LLMs to follow a Chain-of-Thought (CoT) [19] strategy for concentrated vulnerability detection in key code snippets prone to software vulnerabilities. Furthermore, we address the evaluation of generated vulnerability explanations [20] by proposing new evaluation metrics and an automated review method using LLMs.
@@
Through a comprehensive evaluation of two widely used vulnerability datasets (i.e., SeVC [21] and DiverseVul [11]), we find that LLMVulExp can enable the LLMs to detect vulnerabilities with a high accuracy (e.g., with an F1 score over 90% on SeVC) and provide an effective explanation. We also find that the key code extraction following a CoT strategy can improve the performance of vulnerability detection by a large margin.
@@
Paper Organization. Section II summarizes the related work. Section III presents the methodology of our study. Section IV discusses the results to our research questions. Section V discusses the implications of our study. Section VI discusses the threats to validity. Section VII concludes the paper.
II Related Work
@@
II-A Vulnerability Detection
Prior studies proposed a series of deep learning approaches [1, 2, 8, 9, 10] to detect vulnerabilities. These studies utilized labeled data with and without vulnerability to train neural networks and capture their semantic characteristics.
Apart from deep learning approaches, recent advancements in LLMs have significantly influenced the field of software vulnerability detection. Advanced methodologies such as zero-shot prompting [12, 22], in-context learning [23, 24], and fine-tuning [16, 15, 11] have been employed in vulnerability detection. Cheshkov et al. [20] evaluated the performance of the ChaptGPT and GPT-3 on vulnerability detection tasks and found they failed to classify vulnerable codes in binary and multi-label settings. Gao et al. [25] proposed a benchmark for using LLMs in vulnerability detection and validated that on simple datasets with few-shot prompting, LLMs can achieve performance that is comparable to or exceeds that of deep learning methods. Nong et al. [26] focused on evaluating the performance improvement of vulnerability detection tasks using chain-of-thought prompting. By introducing chain-of-thought forms based on the semantic structure of code, LLMs can achieve higher detection accuracy. Sun et al. [27] further assessed the true reasoning capabilities of LLMs by decoupling their vulnerability reasoning abilities. They found that supplementing LLMs with high-quality vulnerability-related knowledge and contextual information can enhance their performance. Yusuf et al. [28] discovered through experiments that natural language instructions enhance the performance of vulnerability detection tasks across multiple programming languages. Steenhoek et al. [29] surveyed and evaluated eleven LLMs for their capabilities of vulnerability detection, utilizing various types of prompts, including in-context learning and chain-of-thought, which indicates the accuracy limitation of directly applying LLMs to vulnerability detection without finetuning.
@@ II-B Vulnerability Explanation
The VulDeeLocator  [30] enhances a Bi-LSTM detector by incorporating an inner multiplication layer, which aids in forecasting vulnerable statements based on the outputs of this layer. In a similar vein, both IVDetect [2] and LineVul [31] develop vulnerability detection models that utilize subgraphs or attention weights derived from the trained detectors to identify vulnerable statements. VELVET [32] prioritizes vulnerable statements by integrating graph-based neural networks with sequence-based neural networks. LineVD [33] addresses statement-level vulnerability detection as a node classification challenge, employing graph neural networks combined with a transformer-based model on PDG for comprehensive learning. VulTeller [34] emphasizes control flows and taint flows to detect more accurate dependencies for localizing vulnerabilities. VulExplainer  [35] is proposed to locate the fine-grained information pertaining to the vulnerability based on graph neural networks. Coca [36] is a dual-perspective contrastive learning enhancement strategy to improve vulnerability explanation methods based on graph neural networks.
@@
III Methodology
Figure 2: Overview of our study.
To address the current gap in generative vulnerability explanation models and enhance the ability of LLMs to detect and analyze software vulnerabilities, we propose a comprehensive framework for fine-tuning and evaluating specialized models for both vulnerability detection and explanation. Figure 2 presents an overview of our framework, namely, LLMVulExp. Specifically, our framework consists of four core stages: ❶ open-source vulnerability data collection, ❷ automated vulnerability explanation annotation based on prompt engineering, ❸ specialized fine-tuning of vulnerability detection and explanation through instruction-based fine-tuning, and ❹ evaluation of generative vulnerability explanation capabilities.
❶ Fine-tuning and Evaluation Vulnerability Dataset Collection: Enhancing the specialized capabilities of LLMs relies on large quantities of high-quality domain-specific data. In the context of vulnerability detection and explanation, the authenticity of the vulnerability code, the diversity of vulnerability types, and the sufficiency of examples of each type are particularly important. In this paper, we conduct the study on two datasets: (1) SeVC [21], which contains four core types and over 50,000 vulnerable code snippets, and (2) DiverseVul [11], which covers 295 real open-source projects and 150 CWE types.
TABLE I: Statistics of the studied datasets.
@@ III Methodology
Semantics-based Vulnerability Candidate (SeVC) dataset includes 126 distinct Common Weakness Enumeration (CWE) types of vulnerabilities with 56,395 vulnerable samples and 364,232 non-vulnerable ones. The SeVC dataset is categorized into four primary groups based on the underlying causes of the vulnerabilities: Library/API Function Call, Array Usage, Pointer Usage and Arithmetic Expression. The SeVC dataset includes a significant number of vulnerabilities for each of its four types, making it suitable for fine-tuning and evaluating models designed to detect and explain specific types of vulnerabilities.
@@
DiverseVul is a C/C++ vulnerable source code dataset, which includes 18,945 vulnerable functions and 330,492 non-vulnerable functions derived from 7,514 commits, encompassing 150 CWEs. DiverseVul is currently the largest real-world C/C++ vulnerability dataset, characterized by longer code snippets, coverage of more diverse projects and CWE types, and lower label noise. Therefore, it is used to evaluate the capabilities of LLMs in handling more challenging real-world scenarios involving a broader range of vulnerability types. We select the top ten most frequent CWE types to construct a fine-tuning dataset, ensuring a sufficient number of samples and a well-defined number of classes for our multi-class detection task.
@@
We deduplicate the vulnerability samples using a hash method. Then, we downsample the non-vulnerability samples in a 1:1 ratio to the vulnerability samples to obtain a balanced dataset, aiming to reduce training overhead and avoid model bias. The dataset details are shown in Table 1. We split the processed dataset into training, validation, and test sets in an 80%: 10%: 10% ratio for both vulnerability and non-vulnerability samples to conduct our experiments.
@@
❷ Automated Vulnerability Explanation Annotation: Current open-source vulnerability datasets predominantly encompass information such as source code, vulnerability labels, CWE types, and commit messages. However, they lack detailed explanations of the vulnerability logic within the source code, presenting a significant challenge for vulnerability detection techniques to provide corresponding explanations for the detection results. Manually annotating real-world vulnerable code explanations requires extensive software development experience and in-depth knowledge of software vulnerabilities, which incurs high labor and time costs.
@@
To address this challenge, we propose an automated vulnerability explanation annotation method based on prompt engineering using LLMs. This method leverages the contextual learning and instruction-following capabilities of LLMs, utilizing prompt engineering to achieve large-scale, high-quality automated vulnerability explanation annotation. The prompts decompose the explanation goal into three sub-goals: vulnerability discrimination, code location, and specific explanation. By combining instruction-based prompt templates with well-annotated examples, we stimulate the model’s contextual learning capabilities, ensuring the effectiveness of vulnerability explanation annotation.
@@
In this paper, we use GPT-3.5 [37], accessed via the API provided by OpenAI [38], to implement the annotation process. To address our research questions and experimental needs, we annotated 40,491 and 9,161 vulnerability explanation data points across two datasets, respectively. This effort fills the current gap in vulnerability explanation data.
@@
❸ Specialized Fine-Tuning of Vulnerability Detection and Explanation: The automated explanation annotation of open-source vulnerability data creates a large-scale dataset for vulnerability detection and explanation, addressing data bottleneck issues in the fine-tuning process. To enhance the vulnerability detection and explanation capabilities of LLMs (especially open-source models with lower computational overhead), we fine-tune the general LLMs to enable them to detect and explain specific types of vulnerabilities in real code. We use instruction-based prompts to guide tasks, helping LLMs correctly understand task goals and generate standardized outputs. To minimize the computational overhead of the fine-tuning process, we adopt the parameter-efficient fine-tuning technique LoRA [16], significantly reducing time and space costs.
@@
❹ Evaluation of Generative Vulnerability Explanation: The current lack of research on model-generated vulnerability explanations highlights the need for effective evaluation methods for LLM-generated explanations. Similar to the challenges in annotation, manual evaluation demands significant human and time resources. Additionally, the effectiveness of vulnerability explanations must be assessed across multiple dimensions to ensure they provide valuable assistance to developers in real-world scenarios. To efficiently evaluate vulnerability explanations, we propose an evaluation method based on three dimensions: accuracy, clarity, and actionability. By leveraging prompt engineering, we develop an automated LLM evaluation method. Expert manual verification is employed to validate both the quality of the specialized vulnerability explanation model’s outputs and the feasibility of the LLM-based automated evaluation scheme.
III-A Vulnerability Interpretation Enhancement Prompting